### v3

# GNN-ACLP: Graph Neural Networks Based Analog Circuit Link Prediction

Abstract
Circuit link prediction identifying missing component connections from incomplete netlists is crucial in automating analog circuit design. However, existing methods face three main challenges: 1) Insufficient use of topological patterns in circuit graphs reduces prediction accuracy; 2) Data scarcity due to the complexity of annotations hinders model generalization; 3) Limited adaptability to various netlist formats. We propose GNN-ACLP, a Graph Neural Networks (GNNs) based framework featuring three innovations to tackle these challenges. First, we introduce the SEAL (Subgraphs, Embeddings, and Attributes for Link Prediction) framework and achieve port-level accuracy in circuit link prediction. Second, we propose Netlist Babel Fish, a netlist format conversion tool leveraging retrieval-augmented generation (RAG) with a large language model (LLM) to enhance the compatibility of netlist formats. Finally, we construct SpiceNetlist, a comprehensive dataset that contains 775 annotated circuits across 10 different component classes. Experiments demonstrate accuracy improvements of 16.08% on SpiceNetlist, 11.38% on Image2Net, and 16.01% on Masala-CHAI compared to the baseline in intra-dataset evaluation, while maintaining accuracy from 92.05% to 99.07% in cross-dataset evaluation, exhibiting robust feature transfer capabilities.
keywords:
Circuit Link Prediction , SEAL , GNNs , LLM , RAG , EDA
journal: Knowledge-Based Systems
[…]
[itmo]organization=HDU-ITMO Joint Institute, Hangzhou Dianzi University, addressline=No. 1158, 2nd Avenue, Xiasha Higher Education Zone, Jianggan District, city=Hangzhou, postcode=310018, state=Zhejiang Province, country=China
[…]
[llp]organization=Intelligent Information Processing Laboratory, Hangzhou Dianzi University, addressline=No. 1158, 2nd Avenue, Xiasha Higher Education Zone, Jianggan District, city=Hangzhou, postcode=310018, state=Zhejiang Province, country=China
[…]
[zju]organization=College of Media Engineering, Communication University of Zhejiang, addressline=No. 998, Xueyuan Street, Xiasha Higher Education Zone, Jianggan District, city=Hangzhou, postcode=310018, state=Zhejiang Province, country=China
[…]
[scbc]organization=School of Computer Science and Technology, South China Business College, Guangdong University of Foreign Studies, addressline=181 Liangtian Middle Road, Baiyun District, city=Guangzhou, postcode=510545, state=Guangdong Province, country=China
[…]
[cam]organization=Department of Computer Science and Technology, University of Cambridge, addressline=William Gates Building, 15 JJ Thomson Avenue, city=Cambridge, postcode=CB3 0FD, state=Cambridgeshire, England, country=United Kingdom
1 Introduction
Analog circuit design has mainly been performed manually by experienced engineers. However, analog design is susceptible to layout geometry [1] and even experienced engineers face difficulties when designing them. Therefore, automating analog circuit design has gained significant attention recently [2]. One main task of analog design automation is circuit link prediction, which infers missing component interconnections from incomplete netlists.
There are three conventional methods for general link prediction: heuristic methods, latent feature methods and learning-based methods.
Heuristic methods can be divided into two categories: local similarity metrics, which include the Jaccard index [3, 4], common neighbors [5], preferential attachment [6], and resource allocation [7], and global similarity metrics, including the Katz index [8] and SimRank [9]. Although heuristic methods offer simplicity and interpretability through predefined topological features, their generalizability is limited by dependence on predefined similarity metrics that encode fixed assumptions about link formation, leading to their incompetence to adapt to the structural heterogeneity of diverse networks [10, 11]. Therefore, they are not commonly used in circuit link prediction.
Latent-feature methods compute latent properties or node representations, typically through factorizing a specific matrix derived from the network. One of the classic latent-features methods is matrix factorization [12], which decomposes the observed adjacency matrix into a low-rank latent-embedding matrix and its transpose. Another classic kind of latent-feature methods are network embedding methods that learn low-dimensional node representations, for instance, DeepWalk [13] and graph2vec [14]. While effective, the generalization capability of latent-feature methods is often constrained by their reliance on specific factorization assumptions and techniques [11].
In the Electronic Design Automation (EDA) field, where circuit netlists represent graph-structured data, machine learning (ML) methods have gained more prominence. Genssler et al. proposed a novel method using brain-inspired hyperdimensional computing (HDC) for encoding and recognizing gate-level netlists [15]. However, the high-dimensional vector operations involved in HDC led to significant computational and storage overhead, especially when dealing with large-scale graph data [16]. Luo et al. proposed a novel neural model called Directional Equivariant Hypergraph Neural Network (DE-HNN) for effective representation learning on directed hypergraphs, particularly for circuit netlists in chip design [17]. However, DE-HNN’s hypergraph diffusion and tensor operations may result in exponentially increasing computational costs as the hyperedge order increases, making it challenging to scale for large chip design [18].
GNNs can leverage the inherent graph structure of circuit netlists, making them well-suited for circuit link prediction. However, a critical challenge is the lack of annotated training data due to high human workload and data sensitivity [19]. Furthermore, the diversity of netlist formats poses significant challenges in analog design automation. Additionally, customized netlist formats are often created by modifying standard formats. Consequently, different datasets are often created in different formats, complicating the training processes for machine learning methods.
To address these challenges, we propose GNN-ACLP, a GNN-based method for analog circuit link prediction, as illustrated in Figure 1. Additionally, we introduce SpiceNetlist, a large-scale netlist dataset with statistics presented in Table 1. 1) We formulate circuit link prediction as an undirected graph link prediction problem, where nodes represent circuit components and edges denote connections. To solve this task, we leverage the SEAL framework, trained using DRNL one-hot encoding augmented with original component node features for enhanced representation learning. 2) We introduce Netlist Babel Fish, a framework featuring RAG for LLM-based interpretation and enabling bidirectional netlist format conversion through domain-specific knowledge integration, to overcome the challenge of differing dataset formats. 3) Finally, we develop a robust preprocessing pipeline to detect and correct content errors and syntactic inconsistencies across SPICE netlist datasets. We then apply the pipeline to two small-scale datasets and integrate them into a single, unified large-scale dataset, SpiceNetlist, to support the training and evaluation of circuit link prediction methods. SpiceNetlist has 775 annotated circuits in both SPICE and JSON formats, featuring 10 component types. The experimental results demonstrate significant performance gains in intra-dataset evaluation, achieving accuracy improvements of 16.08% on SpiceNetlist, 11.38% on Image2Net, and 16.01% on Masala-CHAI compared to the baseline method. Furthermore, our method maintains accuracy from 92.05% to 99.07% in cross-dataset evaluation, exhibiting robust feature transfer capabilities.
The contributions of our work are outlined below:
  • 1.
    GNN-ACLP, a port-level accurate circuit link prediction method based on the SEAL framework using graph neural networks (GNNs). To the best of our knowledge, this is the first work to address circuit link prediction at the port level, as opposed to the conventional component-level methods.
  • 2.
    Netlist Babel Fish, a netlist format converter enabling port-level netlist parsing across multiple formats, leveraging RAG for LLM-based interpretation.
  • 3.
    SpiceNetlist, a novel circuit netlist dataset comprising 775 annotated circuits in both SPICE and JSON formats, featuring 10 component types to facilitate the training and evaluation of circuit link prediction methods.
Figure 1: Illustration of GNN-ACLP. We first preprocess SPICE netlist datasets by correcting errors and syntactic inconsistencies, generating faultless JSON netlists. We then enable bidirectional format conversion via Netlist Babel Fish, a RAG-based LLM framework plus domain-specific knowledge. The netlists are then modeled as graphs, and we apply SEAL for circuit link prediction.
Table 1: SpiceNetlist statistics.
Dataset Graphs Classes Avg. Nodes Avg. Edges
SpiceNetlist 775 10 13.78 15.60
2 Preliminaries and problem formulation
The circuit link prediction problem is a part of the circuit completion problem. One can define the circuit completion problem as follows:
Problem 2.1 (circuit completion problem) Let $v_{0}$ and $v_{1}$ represent the netlists of two circuit schematics, where $v_{1}$ is a partial copy of $v_{0}$ missing all connections associated with a component $x$ that are present in $v_{0}$. The circuit completion problem involves predicting the type of the missing component $x$ and all of its missing connections, where $x\in\{1,2,3,\ldots,k\}$.
Here, netlists define the connections between circuit components, including ports, each with a unique ID and type [20].
There are several challenges in solving circuit completion problems. First, the problem implicitly assumes an explicit criterion for determining a circuit’s validity, which is necessary for verifying a solution’s correctness. Although some checks can be performed, no definitive criteria exist to validate a circuit netlist. Second, it may not be relevant to one’s interests even if a circuit is valid. As long as both $v_{0}$ and $v^{\prime}_{0}$ are netlists of designs that are considered ”interesting” (where ”interesting” is defined by the specific application), it may be acceptable to predict either as a valid completion of the partial netlist $v_{1}$. However, $v^{\prime}_{0}$ may be a trivial completion irrelevant to the particular application being studied [21, 22]. To address these challenges, we discard any auxiliary information from the netlists and convert each netlist into an undirected graph, denoted as $G=(V,E,\phi)$. In this graph, the set of vertices $V$ represents the ports of all components in the netlist, while the edges denote the connections between these ports in $E$ that link the corresponding vertices. Additionally, each vertex $v$ has an integer type, denoted as $\phi(v)$, taken from the set $\{1,2,3,\ldots,k\}$.
We then define the circuit link prediction problem as follows:
Problem 2.2 (circuit link completion problem on graphs) Let $G=(V,E,\phi)$ be a graph of a netlist with $\phi(v)\in[k]$ where $[k]:=\{1,2,3,\ldots,k\}$. Let $G^{\prime}$ be a graph obtained by removing an arbitrary vertex $u\in V$ from $G$. Given the value of $\phi(u)$ and $G^{\prime}$, compute the set of neighbors of $u$ in $G$.
Here, the set of neighbors of a node u is defined as the following set: $\{v\in V:(u,v)\in E\}$. We propose data-driven solutions to tackle the problem. Specifically, we learn a map, ${\hat{\xi}}:{G},[k]\rightarrow 2^{V}$, that takes as input a graph $G$ and a component type $\phi\in[k]$. This mapping returns a subset of vertices (the power set of vertices $V$, represented as $2^{V}$) that indicate the connections for the ports of components within graph $G$. We provide visual illustrations of this problem in Figure 2-3.
Figure 2: Illustration of the circuit link prediction problem. a) Schematic representation of a netlist, b) Schematic representation of a netlist with a missing component.
Figure 3: Architecture of our circuit link prediction architecture with SEAL. Netlists are converted into graph representations and processed by SEAL, which extracts and encloses subgraphs to generate training samples for GNN-based link prediction. We implement DGCNN as the GNN architecture.
3 Link prediction through GNNs
In Problem 2.2, we are given a graph $G$ containing various node types and a distinguished node $v$. The objective of the problem is to predict the neighborhood of $v$ within $G$. This problem can be simplified by determining whether an edge should exist between $v$ and each neighboring node $u$ in $G$. The goal is to assess whether two nodes in the graph will likely be connected by an edge based on the other nodes’ connectivity. Therefore, this connection test can be framed as a link prediction problem.
We select the SEAL framework [23], which we illustrate in Figure 3. The idea is to enclose subgraphs around the links in a graph representing positive class instances to extract training data. For a pair of nodes $(x,y)$, an enclosing subgraph is defined as a subgraph that includes the h-hop neighborhood of both $x$ and $y$. SEAL also utilizes a node labeling approach denoted as Double-Radius Node Labeling (DRNL) to label nodes within the subgraphs. DRNL aims to identify the different roles of nodes while preserving structural information. We adopt the SEAL framework for link prediction in circuit graphs and train it using DRNL one-hot encoding combined with the original component node features.
It is important to note that our work may not address the circuit link prediction problem on graphs since there can be multiple valid ways to link ports in a partial circuit [24, 25, 26]. Therefore, the proposed method may occasionally encounter failures. However, we show through extensive experiments that graph machine-learning methods yield reliable results on real-world datasets to support human experts in the design, synthesis, and evaluation processes.
4 Netlist Babel Fish: Netlist Format Converter
To address the challenge of differing dataset formats, we develop Netlist Babel Fish, a framework that enables bidirectional conversion between netlist formats through integration with domain-specific information. Netlist Babel Fish integrates an LLM with a RAG system grounded in a structured SPICE knowledge base. This knowledge base includes brief references of the SPICE devices and statements [27], methodologies for the netlist process, supported component libraries with port assignment specifications, and examples illustrating practical implementations. We employ DeepSeek-V2.5 [28, 29], DeepSeek-V3-0324 [30, 31] and Qwen3-30B-A3B [32] as the LLM in this framework. We show the workflow of Netlist Babel Fish converting a SPICE netlist to a custom JSON format [33] in Figure 4. This workflow is also reversible for conversion from the custom JSON format to SPICE.
Figure 4: Workflow of Netlist Babel Fish. A SPICE netlist is first processed by an embedded query system that retrieves relevant information from our circuit knowledge base, which includes netlist preprocessing methodologies, supported component libraries including port assignment specifications, and exemplar netlists in both SPICE and JSON formats. The netlist and retrieved data are then fed into an LLM, which generates the corresponding JSON netlist as output.
The current implementation focuses on SPICE and specific JSON format conversions. However, its modular architecture can theoretically support arbitrary netlist format transformations with appropriate domain knowledge bases. The system’s extensibility stems from its decoupled knowledge representation framework, where format-specific conversion rules are explicitly encoded.
5 SpiceNetlist: Datasets Rectification
For the circuit-link-prediction problem, few datasets are available. Moreover, existing SPICE netlist datasets often contain content errors [33] and syntactic inconsistencies, leading to their incompatibility with Netlist Babel Fish for proper parsing.
To address this issue, we utilize PySpice [34] to automatically detect and filter out SPICE netlists with syntactic inconsistencies. Subsequently, we convert all datasets to JSON format with Netlist Babel Fish, and perform manual verification to identify and rectify netlists with content errors in both formats. We illustrate this preprocessing pipeline in Figure 5.
Following this pipeline, we integrate two datasets — Kicad Github [35] and AMSNet [36] as one unified large-scale dataset, SpiceNetlist. Table 1 shows SpiceNetlist’s statistics. We also apply the preprocessing pipeline to three additional datasets: Image2Net [^1] Image2Net is provided by Yiren Pan (panyiren@hdu.edu.cn), who is currently writing a paper about his findings, and originated from the 2024 China Postgraduate IC Innovation Competition - EDA Elite Challenge Contest [37] Image Dataset., Masala-CHAI [38] and AnalogGenie [39].
Figure 5: SPICE netlist datasets preprocessing pipeline. SPICE netlists are first validated using PySpice, with syntactically inconsistent netlists being filtered out. The remaining netlists undergo conversion to JSON format via Netlist Babel Fish, where content errors are manually corrected. The resulting faultless JSON netlists are then processed back through Netlist Babel Fish to generate faultless SPICE equivalents.
6 Experiments
6.1 Dataset Processing
In our experiments, we utilize SpiceNetlist, Image2Net, Masala-CHAI, and AnalogGenie [^2] Given the substantial size of AnalogGenie, we randomly select a representative 10% subset for our experiments.. Table 2 shows the statistics of all datasets, and Table 3 lists all components featured by all datasets.
Table 2: Statistics of all datasets utilized in our experiments.
Dataset Graphs Classes Avg. Nodes Avg. Edges
SpiceNetlist 775 10 13.78 15.60
Image2Net 1960 13 26.36 37.80
Masala-CHAI 4057 13 15.72 19.97
AnalogGenie 4064 9 63.89 168.73
Table 3: Component types and their labels/ports in the datasets.
Component Label Ports
PMOS/NMOS PMOS/NMOS Drain, Source, Gate
Voltage Source Voltage Pos, Neg
Current Source Current In, Out
BJT (NPN/PNP) NPN/NPN_cross/PNP/PNP_cross Base, Emitter, Collector
Diode Diode In, Out
DISO Amplifier Diso_amp InN, InP, Out
SISO Amplifier Siso_amp In, Out
DIDO Amplifier Dido_amp InN, InP, OutN, OutP
Passive Components (Cap, Ind, Res) Cap, Ind, Res Pos, Neg
We opt for an approach similar to batching in the graph learning domain, where we stack the adjacency matrices in diagonal blocks to represent the entire dataset as a single huge graph. Let ${{A}}_{i}\in[0,1]^{N_{i}\times N_{i}}$ represent the adjacency matrix of graph $G_{i}$, where $N_{i}$ is the number of nodes. Then, the resulting graph will have $N:=\sum_{1\leq j\leq n}N_{j}$ nodes. The stacked adjacency matrix ${A}$ with dimensions, $N\times N$, and the corresponding vector with concatenated node features are defined as:
${A}=\begin{bmatrix}{A}_{1}&&\\ &\ddots&\\ &&{A}_{n}\end{bmatrix},\quad{X}=\begin{bmatrix}{X}_{1}\\ \vdots\\ {X}_{n}\end{bmatrix}$
Link prediction methods typically require a single large graph as input to learn the network structure [40, 41]. Therefore, stacking all graphs in the training dataset allows us to utilize off-the-shelf GNN methods for prediction with minimal additional computational overhead. The large adjacency matrix can also be stored in a compressed format using sparse data structures.
6.2 Experimental Setup
We conduct two distinct evaluation paradigms: intra-dataset evaluation and cross-dataset evaluation. For intra-dataset evaluation, models are trained and tested on the same dataset, specifically SpiceNetlist, Image2Net and Masala-CHAI [38]. For cross-dataset evaluation to assess the generalization capability of our method, we first examine models trained on one dataset and evaluated on another from these three datasets, followed by additional validation using AnalogGenie [39].
For both evaluation paradigms, we employ two experimental configurations: one using the conventional dataset splitting strategy (train, validation, and test splits), and another employing the 5-fold cross-validation strategy, given the prevalent challenge of limited sizes in existing datasets. Here, the traditional split serves as an ablation study to assess the impact of 5-fold cross-validation, ensuring robustness in performance evaluation. For both configurations, we maintain a fixed data split ratio of 70% training, 20% test, and 10% validation. We also conduct experiments with [35] as the baseline, which uses the conventional dataset splitting strategy.
We adopt the SEAL framework with the following Pytorch Geometric implementation: batch size = $1$, $2$-hop neighborhoods and maximum $50$ training epochs with early stopping. The early stopping criterion only engages once test accuracy shows improvement and exceeds $0.5000$. Thereafter, training terminates if no improvement $\geq 0.0001$ is observed for three consecutive epochs. We use the one-hot encoding of node labels generated through DRNL and concatenate them with the one-hot encoding of the component types. For SpiceNetlist, we set the learning rate to $61e-$ 6. For Image2Net and Masala-CHAI, we set the learning rate to $61e-$ 6 for the baseline and $86e-$ 8 for our method.
The training graph is constructed using the adjacency stacking approach from Section 3. During each experiment, we remove one random vertex from each test graph and predict its connections. We repeat the experiment 10 times and report the average accuracy and ROC-AUC.
6.3 Results and Analysis
We present a comprehensive comparison of circuit link prediction performance between our work and the baseline [35] in Table 4 and Figure 6-8. The experimental results demonstrate substantial performance improvements across all evaluation scenarios:
In the intra-dataset evaluation, 1) We achieve accuracy gains of 11.08% (conventional split) and 16.08% (5-fold cross-validation) on SpiceNetlist; 2) We achieve accuracy gains of 9.18% (conventional split) and 11.38% (5-fold cross-validation) on Image2Net; 3) We achieve accuracy gains of 13.89% (conventional split) and 16.02% (5-fold cross-validation) on Masala-CHAI.
In the cross-dataset evaluation, our work maintains accuracy ranging from 92.05% (trained on Image2Net, tested on SpiceNetlist) to 99.07% (trained on SpiceNetlist, tested on AnalogGenie), demonstrating its strong capability to transfer learned features across different circuit netlist datasets.
Our work achieves promising results, indicating that the proposed framework effectively learns netlist structures and can be utilized for component link prediction tasks.
Table 4: Accuracy and ROC-AUC in intra-dataset evaluation. The bold fonts denote the best and sub-optimal results.
Method SpiceNetlist Image2Net Masala-CHAI
Accuracy ROC-AUC Accuracy ROC-AUC Accuracy ROC-AUC
Baseline 77.96% ± 1.71% 0.8774 ± 0.0056 85.72% ± 1.95% 0.9367 ± 0.0164 80.40% ± 0.92% 0.8995 ± 0.0099
Conv. Split (Ours) 89.04% ± 1.18% 0.9645 ± 0.0097 94.90% ± 0.25% 0.9926 ± 0.0004 94.29% ± 0.28% 0.9892 ± 0.0009
5-fold CV (Ours) 94.04% ± 0.70% 0.9866 ± 0.0019 97.10% ± 0.24% 0.9979 ± 0.0003 96.42% ± 1.24% 0.9959 ± 0.0005
Figure 6: AUC and Accuracy variation tendencies observed in the baseline. Both indicators vibrate violently.
Figure 7: AUC and Accuracy variation tendencies observed in our method with conventional dataset splitting strategy. Both indicators vibrate less violently than the baseline method.
Figure 8: AUC and Accuracy variation tendencies observed in our method with 5-fold cross-validation strategy. Both indicators increase steadily.
Table 5: Accuracy and ROC-AUC in cross-dataset evaluation. The bold fonts denote the best results.
Train Dataset Test Dataset Baseline Performance 5-fold CV (Ours) Performance
Accuracy ROC-AUC Accuracy ROC-AUC
SpiceNetlist Image2Net 82.95% ± 1.45% 0.9163 ± 0.0114 96.94% ± 0.88% 0.9963 ± 0.0014
SpiceNetlist Masala-CHAI 77.00% ± 1.47% 0.8600 ± 0.0030 95.98% ± 1.08% 0.9945 ± 0.0017
Image2Net SpiceNetlist 75.10% ± 3.11% 0.8848 ± 0.0050 92.05% ± 0.75% 0.9867 ± 0.0021
Image2Net Masala-CHAI 75.47% ± 1.28% 0.8706 ± 0.0048 96.12% ± 0.66% 0.9962 ± 0.0007
Masala-CHAI SpiceNetlist 72.84% ± 8.50% 0.8678 ± 0.0424 93.06% ± 1.66% 0.9859 ± 0.0019
Masala-CHAI Image2Net 82.35% ± 8.01% 0.9293 ± 0.0151 97.18% ± 0.85% 0.9974 ± 0.0004
SpiceNetlist AnalogGenie 69.99% ± 2.17% 0.7822 ± 0.0330 99.07% ± 0.82% 0.9995 ± 0.0003
Image2Net AnalogGenie 71.23% ± 2.32% 0.7950 ± 0.0402 98.73% ± 1.32% 0.9997 ± 0.0002
Masala-CHAI AnalogGenie 74.19% ± 2.58% 0.8107 ± 0.0378 98.68% ± 1.07% 0.9997 ± 0.0002
6.4 Ablation Study
To demonstrate the utility of the SEAL framework underlying our method, we conduct an ablation study by replacing the SEAL framework with several alternative methods, including heuristic methods, latent feature methods, and other GNN-based methods, across SpiceNetlist, Image2Net, and Masala-CHAI, using the 5-fold cross-validation strategy. We present the results in Table 6-8333Since some of the frameworks are infeasible due to high computational complexity, we only include the results for the methods which are feasible for the experiment (i.e., OOM indicates the method needs more than 48 GB GPU RAM, and OOT indicates the method takes more than 48 hours)..
Framework SpiceNetlist Image2Net Masala-CHAI
Accuracy ROC-AUC Accuracy ROC-AUC Accuracy ROC-AUC
common neighbors [5] 75.86% ± 0.40% 0.7586 ± 0.0040 88.36% ± 0.15% 0.8836 ± 0.0015 86.16% ± 0.16% 0.8616 ± 0.0016
Jaccard [4] 74.92% ± 0.89% 0.7586 ± 0.0040 88.34% ± 0.16% 0.8836 ± 0.0015 86.14% ± 0.15% 0.8616 ± 0.0016
preferential attachment [6] 59.16% ± 1.74% 0.6198 ± 0.0059 65.33% ± 0.42% 0.6972 ± 0.0023 65.65% ± 0.11% 0.7043 ± 0.0023
Adamic-Adar [42] 75.82% ± 0.41% 0.7586 ± 0.0040 88.35% ± 0.15% 0.8836 ± 0.0015 86.16% ± 0.15% 0.8616 ± 0.0016
resource allocation [7] 75.82% ± 0.41% 0.7586 ± 0.0040 88.35% ± 0.15% 0.8836 ± 0.0015 86.16% ± 0.15% 0.8616 ± 0.0016
Katz [8] 85.08% ± 0.32% 0.8515 ± 0.0033 OOT OOT OOT OOT
PageRank [43] 59.79% ± 1.47% 0.6277 ± 0.0059 OOT OOT OOT OOT
SimRank [9] 79.47% ± 0.16% 0.7952 ± 0.0015 90.95% ± 0.03% 0.9096 ± 0.0004 OOM OOM
WLNM [44] 69.33% ± 0.36% 0.8170 ± 0.0025 89.15% ± 0.03% 0.9779 ± 0.0002 91.22% ± 0.04% 0.9757 ± 0.0001
WEGL [45] 64.17% ± 3.26% 0.6036 ± 0.0384 80.92% ± 2.46% 0.8230 ± 0.0207 77.53% ± 4.71% 0.7752 ± 0.0391
GNN-ACLP(SEAL) 94.04% ± 0.70% 0.9866 ± 0.0019 97.10% ± 0.24% 0.9979 ± 0.0003 96.42% ± 1.24% 0.9959 ± 0.0005
Table 6: Ablation study comparison results of SEAL and heuristic methods. The bold fonts denote the best and sub-optimal results.
Framework SpiceNetlist Image2Net Masala-CHAI
Accuracy ROC-AUC Accuracy ROC-AUC Accuracy ROC-AUC
SBM [46] 58.57% ± 1.75% 0.6109 ± 0.0031 59.77% ± 0.11% 0.6032 ± 0.0015 60.91% ± 0.07% 0.6144 ± 0.0013
BMF [47] 52.26% ± 0.19% 0.5405 ± 0.0031 58.96% ± 0.09% 0.6760 ± 0.0008 58.23% ± 0.08% 0.6621 ± 0.0010
graph2vec [14] 65.75% ± 2.74% 0.7315 ± 0.0283 60.69% ± 0.41% 0.6938 ± 0.0089 OOT OOT
AROPE [48] 71.56% ± 0.15% 0.8048 ± 0.0027 71.50% ± 0.08% 0.8041 ± 0.0010 71.52% ± 0.14% 0.8044 ± 0.0022
SPC [49] 71.53% ± 3.26% 0.8197 ± 0.0108 OOT OOT OOT OOT
GNN-ACLP(SEAL) 94.04% ± 0.70% 0.9866 ± 0.0019 97.10% ± 0.24% 0.9979 ± 0.0003 96.42% ± 1.24% 0.9959 ± 0.0005
Table 7: Ablation study comparison results of SEAL and latent feature methods. The bold fonts denote the best and sub-optimal results.
Framework SpiceNetlist Image2Net Masala-CHAI
Accuracy ROC-AUC Accuracy ROC-AUC Accuracy ROC-AUC
VGAE [50] 63.80% ± 2.00% 0.7197 ± 0.0099 71.58% ± 1.18% 0.8318 ± 0.0078 69.37% ± 1.54% 0.8115 ± 0.0114
GraphSAGE [51] 65.07% ± 2.94% 0.7683 ± 0.0172 78.45% ± 0.26% 0.9147 ± 0.0025 77.74% ± 0.19% 0.9108 ± 0.0009
GCN+LRGA [52] 50.89% ± 1.21% 0.7029 ± 0.0268 OOT OOT OOT OOT
GNN-ACLP 94.04% ± 0.70% 0.9866 ± 0.0019 97.10% ± 0.24% 0.9979 ± 0.0003 96.42% ± 1.24% 0.9959 ± 0.0005
Table 8: Ablation study comparison results of SEAL and other GNN-based methods. The bold fonts denote the best and sub-optimal results.
We believe that the superior performance of SEAL can be attributed to its inherent adaptability for analog circuit link prediction, wherein the underlying graphs exhibit both sparsity and large-scale topology.
Heuristic methods, which rely on closed-triangle assumptions or preferential attachment mechanisms, prove inadequate in sparse graphs due to their inability to capture structural idiosyncrasies [53]. Latent feature methods fails either since insufficient node interactions prevent meaningful learning of high-order correlations needed for latent semantic extraction [54]. They may also erroneously encode unobserved links as zero-weight edges, creating inconsistency between partial and complete graph representations [55]. As for other GNN-based methods here, they are node-based and compute two node representations independently of each other without considering their relative positions and associations. SEAL, on the other hand, is a subgraph-based method and extracts an enclosing subgraph around each target link. Therefore, it better models the associations between two target nodes in a sparse graph. The node labeling step in SEAL also helps model the associations between the two target nodes, especially when the graph is sparse [56].
The computational resources required by different frameworks for predicting on large circuit graphs are also a critical consideration. Katz [57], PageRank [58], SimRank [59], graph2vec [60], SPC [61], GCN [62] and VGAE [^4] VGAE faces a trade-off in node feature decomposition: a large latent dimension increases computational overhead, while a small one may fail to capture essential latent factors [63]. Additionally, VGAE-based methods often suffer from training instability, and correcting this deviation introduces additional computational costs. [64, 65] suffer from high computational complexity, making them inefficient for large-scale circuit graphs. Therefore, when applied to circuit graphs that are large, these methods exhibit prohibitive resource demands and may even OOM or OOT. For instance, SimRank consumes nearly 40 GB of GPU memory on Image2Net, graph2vec requires more than 20 hours for training on SpiceNetlist, and VGAE occupies 34 GB of memory on Masala-CHAI. In contrast, SEAL completes training on each individual dataset (SpiceNetlist, Image2Net, and Masala-CHAI) in under 10 hours with significantly lower GPU memory.
7 Discussion
While our GNN-ACLP showcases significant promise in circuit link prediction, limitations exist. To comprehensively analyzing the limits of GNN-ACLP, failed examples are shown in Table 9-1100footnotemark: 0. Specifically, we train and test GNN-ACLP and other methods on AnalogGenie with the 5-fold cross-validation strategy.
Framework AnalogGenie
Accuracy ROC-AUC
common neighbors 95.32% ± 0.03% 0.9532 ± 0.0003
Jaccard 95.32% ± 0.03% 0.9532 ± 0.0003
preferential attachment 69.64% ± 0.05% 0.7403 ± 0.0006
Adamic-Adar 95.32% ± 0.03% 0.9532 ± 0.0003
resource allocation 95.32% ± 0.03% 0.9532 ± 0.0003
Katz OOT OOT
PageRank OOT OOT
SimRank OOM OOM
WLNM 97.46% ± 0.04% 0.9978 ± 0.0000
WEGL 82.69% ± 3.14% 0.8778 ± 0.0295
GNN-ACLP(SEAL) OOT OOT
Table 9: Performance comparison of GNN-ACLP and heuristic methods on AnalogGenie. The bold fonts denote the best and sub-optimal results.
Framework AnalogGenie
Accuracy ROC-AUC
SBM 56.05% ± 0.02% 0.5629 ± 0.0004
BMF 64.38% ± 0.03% 0.7776 ± 0.0002
graph2vec OOT OOT
AROPE 71.55% ± 0.20% 0.8046 ± 0.0027
SPC OOT OOT
GNN-ACLP(SEAL) OOT OOT
Table 10: Performance comparison of GNN-ACLP and latent feature methods on AnalogGenie. The bold fonts denote the best and sub-optimal results.
Framework AnalogGenie
Accuracy ROC-AUC
VGAE OOM OOM
GraphSAGE 79.36% ± 0.17% 0.9296 ± 0.0020
GCN+LRGA OOT OOT
GNN-ACLP(SEAL) OOT OOT
Table 11: Performance comparison of GNN-ACLP and other GNN methods on AnalogGenie. The bold fonts denote the best results.
The observed OOT issue in GNN-ACLP for AnalogGenie may be attributed to the inherent computational complexity of SEAL, particularly due to its node labeling approach [56]. Notably, SEAL’s performance tends to degrade when applied to highly dense graph structures since its labeling approach struggles to distinguish critical circuit subgraphs [66], a key challenge in AnalogGenie’s highly dense networks.
To address these challenges, we will optimize SEAL’s enclosing subgraph extraction by filtering electrically irrelevant nodes and introduce circuit-specific labeling approach to reduce complexity without sacrificing accuracy in the future. Furthermore, we are exploring alternative frameworks capable of more effective structural representation learning in dense graphs, thereby overcoming SEAL’s limitations in such scenarios.
8 Conclusion
We propose GNN-ACLP, a novel graph neural networks based framework with SEAL framework integration and achieve port-level accuracy in analog circuit link prediction. We also propose Netlist Babel Fish, a netlist format conversion tool combining RAG with an LLM. Additionally, we create the SpiceNetlist dataset for training and evaluation of circuit link prediction methods. Results of the experiment show that our method achieves state-of-the-art performance.
CRediT authorship contribution statement
Guanyuan Pan: Investigation, Validation, Software, Writing – original draft, Visualization. Tiansheng Zhou: Data curation, Visualization. Bingtao Ma: Writing – review and editing. Yaqi Wang: Conceptualization, Writing – review and editing. Jianxiang Zhao: Writing – review and editing. Zhi Li: Visualization. Yugui Lin: Validation. Pietro Liò: Writing – review and editing. Shuai Wang: Resources, Supervision, Writing – review and editing.
Data Availability
Code and data will be made available on request.